OpenStudy (anonymous):

Suppose oil spills from a ruptured tanker and spreads in a circular pattern. If the radius of the oil spill increases at a constant rate of 3m/s, how fast is the area of the spill increasing when the radius is 40m? Rate of change of area = m^2/s

OpenStudy (anonymous):

\[A(r)=\pi r^2\\ {dA\over dt}=2\pi r{dr\over dt}\\ \quad=2\pi\times 40{\rm m}\times 3{\rm m/s}\\ \quad=240\pi\, {\rm m^2/s} \]

OpenStudy (anonymous):

so its the same thing as yesterday, just with less information given?

OpenStudy (anonymous):

what yesterday?

OpenStudy (anonymous):

things like this Suppose that x=x(t) and y=y(t) are both functions of t. If x2+y2=25, and dx/dt=2 when x=3 and y=4, what is dy/dt? dy/dt =

OpenStudy (anonymous):

right. parametric equations.

OpenStudy (anonymous):

ok, il see what i can do on these other ones, thank you